OpenStudy (anonymous):

Find the equation of a circle if the end points of the diameter are (-2,3) and (2,3)

OpenStudy (mathmate):

Here, find the centre of the circle by taking the mid-point between the two points: C(x,y)=((-2+2)/2,(3+3)/2)=(0,3) The radius is evidently 2 from sqrt((2-0)^2+(3-3)^2)=2 So the equation for a circle with centre at C(0,3) and radius 2 is (x-0)^2+(y-3)^2=2^2 => x^2+(y-3)^2=4